Science Report Oct 27th

Crew 214 Science Report 27-Oct-2019

Submitted by Science Officer Andrew Wheeler

1. EVA Suit Maintenance: The hab has been set up following arrival. All backpacks have been plugged in and fully charged. Helmets are clean and contained in their cloth bags except for one helmet, for which, the cloth bag is missing. Radios have been tested and are fully charged. Communication earpieces remain to be tested for faulty performance.

2. Green Hab: The green hab has been activated and is awaiting water and additional soil base before sowing of this seasons crops.

3. Astronomy Dome Operations: No astronomical observations have been planned for this crew rotation.

4. Science Dome Operations: The Science Dome has been activated and is awaiting EVAs for sample acquisition and sample testing.

5. RAM Operations: The RAM Dome has been activated and is ready for contingencies.

6. EVA: No preparations or activities have been undertaken to date. The access routes have been identified to allow sample locations to be identified.

7. Closed Loop Food Waste Study: This study has been initiated and is ongoing. Day’s waste: 79g; consisting of 323 calories, 10g fat, 5g saturated fat, 1179mg sodium, 52g carbohydrates, 7g fibre, 6g sugar, 7g protein.

8. Additional activities: Nil

Spacesuit Report – May 16th

Crew 212-LATAM-III

Submitted by Spacesuit Engineer: Camilo Andrés Zorro Mendoza

Spacesuit report

Sol: 11

Sim suit used Engineering EVA: 7, 8, 5 and 4.
used for 1 hour and half

Sim suit: 7
Voltage level before EVA: 12.8 V
Voltage level after EVA:
12.5 V

Sim suit: 4
Voltage level before EVA:
13.1 V
Voltage level after EVA:
12.6 V

Sim suit: 5
Voltage level before EVA: 12.6 V
Voltage level after EVA:
5.6 V (it will be changed)

Sim suit: 8
Voltage level before EVA:
13 V
Voltage level after EVA:
12.8 V

Science Report – May 15th

Crew 212-LATAM-III

Submitted by Spacesuit Engineer: Camilo Andrés Zorro Mendoza

Spacesuit report

Sol: 10

Sim suit used in EVA: 9 and 3.
used for 2 hour

Sim suit: 9
Voltage level before EVA: 12.8 V
Voltage level after EVA:
12.6 V

Sim suit: 3
Voltage level before EVA:
13.2 V
Voltage level after EVA:
12.9 V

General commentaries

– I am running a discharging test for the extra batteries. Until this moment, one battery had gone bad.

Spacesuit Report – May 13th

Crew 212-LATAM-III

Submitted by Spacesuit Engineer: Camilo Andrés Zorro Mendoza

Spacesuit report

Sol: 8

Sim suit used EVA: 1, 8, 3 and 5.
used for 2 and 1/2 hour

Sim suit: 1
Voltage level before EVA: 12.8 V
Voltage level after EVA:
12.7 V

Sim suit: 8
Voltage level before EVA: 12.8 V
Voltage level after EVA:
4.1 V

Sim suit: 3
Voltage level before EVA:
13 V
Voltage level after EVA:
12.8 V

Sim suit: 5
Voltage level before EVA: 13.2 V
Voltage level after EVA:
7.6 V

General commentaries

It is need get 10 new batteries for the radios. It was getting fake charge. 4 radios with good battery

Science Report – May 11th

Crew 212-LATAM-III

Submitted by Spacesuit Engineer: Camilo Andrés Zorro Mendoza

Spacesuit report

Sol: 6

Sim suit used: 3 and 5.
used for 1 hour

Sim suit: 3
Voltage level before EVA: 13.6 V
Voltage level after EVA: 12.9 V

Sim suit: 5
Voltage level before EVA: 13. V
Voltage level after EVA: 12.8 V

General commentaries

The batteries of radios are not working properly. In the following days, I will give you a detail report on the number of batteries that should be changed.

Science Report – May 10th

Crew 212-LATAM-III

Submitted by Spacesuit Engineer: Camilo Andrés Zorro Mendoza

Spacesuit report

Sol: 5

Sim suit used for first EVA: 7, 4 and 10.
used for 1 hour

Sim suit: 4
Voltage level before EVA: 12.6 V
Voltage level after EVA: 8.4 V

Sim suit: 7
Voltage level before EVA: 13.1 V
Voltage level after EVA: 12.7 V

Sim suit: 10
Voltage level before EVA: 13.1 V
Voltage level after EVA: 12.9 V

Sim suit used for second EVA: 3, 5, 8 and 9.
used for 1 1/2 hours

Sim suit: 3
Voltage level before EVA: 13.4 V
Voltage level after EVA: 12.8 V

Sim suit: 5
Voltage level before EVA: 13.4 V
Voltage level after EVA: 12.8 V

Sim suit: 8
Voltage level before EVA: 13.5 V
Voltage level after EVA: 12.3 V

Sim suit: 9
Voltage level before EVA: 13.3 V
Voltage level after EVA: 12.8 V

General comments:
– It seems that the battery of sim suit 4 is draining the battery very fast.
– The battery recharger was found. I will proceed to charge the extra batteries and see if any of them have gone bad

Science Report – May 9th

Crew 212 – LATAM III

Submitted by GreenHab Officer Hector Palomeque

1. Solar Observation/Astrophotography/Stellar Clusters:

• The eyepiece of MUSK Telescope was successfully assembled.

• First picture of the sun with MUSK Solar Telescope (a bit cloudy).

• Deep Sky Robotic Telescope still observing a star cluster (interrupted due to the bad weather).

2. X-5:

• The first flight of 2 on 3 drones in a Martian environment was performed during today’s EVA (SOL 4).

• Manual light off and landing did for checking that Drones are working correctly.

• Drones operation with space suit was tested.

• Waiting for the weather forecast to schedule and perform the next experiment (photos of the MDRS Hab and other areas).

3. VESTA:

• One flight of the Drone-Multicopter were performed around MDRS Hab during today’s EVA (SOL 4)

• Interesting points around MDRS Hab selected.

• Drone successfully connected to the computer for planning an automatic mission.

4. Observation Rover:

• Linux system correctly launched and running in a crew’s Laptop.

• Still needing to download the script from UK for uploading it into the Raspberry Pi Rover board.

5. Space Farming:

• Red-Clay collected for preparing a mix for Martian-like regolith.

• Broken-Autoclave used for boiling in order to disinfect/sterilize the material as much as possible (without problem nor danger).

• Broth culture media for selected Chernobyl and Fukushima bacteria successfully prepared and sterilized (no contamination detected in negative control after 24 hours).

• Chernobyl and Fukushima bacteria correctly inoculated in Nutrient Broth.

• Chernobyl and Fukushima bacteria already showing sights of growth.

6. Keep Talking:

• Crewmembers getting a better performance during the game. Some teams already have reached the Difficult Level.

• The fourth day of playing (corresponding for each SOL).

7. Ethnography study:

• Audio-recordings with the information collected during the observations of the Crew 212 behave already sent to the Colombian Aerospace Research and Develop Group (Colombia).

8. Pulse Oximeter Measurement:

• No update to report.

9. Software for macroscopic characterization:

• Nutrient Agar in Petri dishes prepared using the Broken-Autoclave (all correct).

• MDRS soil bacteria cultivated in Nutrient Agar.

• Waiting for bacterial growth (24 hours).

• EVA scheduled for collecting rhizospheric soil of MDRS plants.

Science Report – May 9th

Spacesuit Engineer: Camilo Andrés Zorro Mendoza

Spacesuit report

Sol: 4

Sim suit used: 3, 4, 8 and 9.
used for 2 hours

Sim suit: 3
Voltage level before EVA: 13.4 V
Voltage level after EVA:
12.8 V

Sim suit: 5
Voltage level before EVA: 13 V
Voltage level after EVA:
12.7 V

Sim suit: 8
Voltage level before EVA: 13.2 V
Voltage level after EVA:
10.5 V

Sim suit: 4
Voltage level before EVA: 12.9 V
Voltage level after EVA:
10.6 V

Science Report – May 8th

Crew 212-LATAM-III

Submitted by Spacesuit Engineer: Camilo Andrés Zorro Mendoza

Spacesuit report

Sol: 3

Sim suit used: 8, 3, 9 and 5

Sim suit: 3
Voltage level before EVA: 12.9 V
Voltage level after EVA:
12.6 V

Sim suit: 5
Voltage level before EVA: 13.7 V
Voltage level after EVA:
12.9 V

Sim suit: 8
Voltage level before EVA: 12.8 V
Voltage level after EVA:
10.5 V

Sim suit: 9
Voltage level before EVA: 13.1 V
Voltage level after EVA:
12.7 V

General comments:

It was founded that four batteries of radios does not get charge. Also one sim suit recharger does not work. The radios was numbered.

Science Report – May 07th

Science Report 7 May 2018

Crew 212 – LATAM III

Submitted by GreenHab Officer Hector Palomeque

1. Solar Observation/Astrophotography/Stellar Clusters: The procedure of turning on the MUSK telescope and the dome has been successfully tested. Observations of a galaxy were made with the deep sky objects robotic telescope.

2. X-5: Drones were disassembling for transportation to MDRS: 2 on 3 drones already assembled. Calibration performed. Manual test of 2 on 3 drones scheduled for SOL-3.

3. VESTA: Performing a repair in the landing gear of the drone. First flight test scheduled for May, 09 (SOL 4).

4. Observation Rover: The EVA for today May, 7 (SOL 2) was canceled due to the weather.. A part of the Rover got missing in transit from Grand Junction to MDRS. Critical Component. Solution needed ASAP: Stablish contact with UK.

5. Space Farming: Experiment in standby due to the lack of an operational autoclave. Solution: Dr. Shannon Rupert has explained how to sterilize using a hotplate, alcohol, and bleach.

6. Keep Talking: First round already performed. Subsequent rounds already scheduled.

7. Ethnography study: First observation started yesterday (SOL-1). Waiting for sending audio-information to receive the feedback from the Colombian team project.

8. Pulse Oximeter Measurement: First measurement made yesterday. Data recorded.

9. Software for macroscopic characterization: Code for detecting counts in progress. There is an issue with detecting the petri dish. Solution: process more photographs needed (soil bacteria growing in plates).